the perimeter of a rectangle is 86 inches. The length exceeds the width by 9 inches. Find the length and the width? Where do I start?
Answer by rfer(16322)

2x+2(x+9)=86
2x+2x+18=86
4x=68
x=17 width
x+9=26 length